40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 16 16:02:33 2021, Koji, Update, LSC, REFL11 demod retrofitting IMG_0545.jpegIMG_0547.jpegD040179-A.pdfIMG_0548.jpegIMG_0550.jpeg
    Reply  Thu Mar 18 07:02:27 2021, Koji, Update, LSC, REFL11 demod retrofitting TFs.pdfPSD.pdf
Message ID: 15923     Entry time: Tue Mar 16 16:02:33 2021     Reply to this: 15936
Author: Koji 
Type: Update 
Category: LSC 
Subject: REFL11 demod retrofitting 

I'm going to remove REFL11 demod for the noise check/circuit improvement.

----

  • The module was removed (~4pm). Upon removal, I had to loosen AS110 LO/I out/Q out. Check the connection and tighten their SMAs upon restoration of REFL11.
  • REFL11 configuration / LO: see below, PD: a short thick SMA cable, I OUT: Whitening CH3, Q OUT: Whitening CH4, I MON daughterboard: CM board IN1 (BNC cable)
  • The LO cable for REFL11 was made of soft coax cable (Belden 9239 Low Noise Coax). The vendor specifies that this cable is for audio signals and NOT recommended for RF purposes [Link to Technical Datasheet (PDF)].
    I'm going to measure the delay of the cable and make a replacement.
  • There is a bunch of PD RF Mon cables connected to many of the demo modules. I suppose that they are connected to the PD calibration system which hasn't been used for 8 years. And the controllers are going to be removed from the rack soon.
    I'm going to remove these cables.

----

First I checked the noise levels and the transfer functions of the daughterboard preamp were checked. The CH-1 of the SR785 seemed funky (I can't comprehensively tell yet how it was), so the measurement maybe unreliable.

For the replacement of AD797, I tested OP27 and TLE2027. TLE2027 is similar to OP27, but slightly faster, less noisy, and better in various aspects.

The replacement of the AD797 and whatever-film resistors with LTE2027 and thin-film Rs were straightforward for the I phase channel, while the stabilization of the Q phase channel was a struggle (no matter I used OP27 or TLE2027). It seems that the 1st stage has some kind of instability and I suffered from 3Hz comb up to ~kHz. But the scope didn't show obvious 3Hz noise.

After a quite bit of struggle, I could tame this strange noise by adjusting the feedback capacitor of the 1st stage. The final transfer functions and the noise levels were measured. (To be analyzed later)

----

Now the REFL11 LO cable was replaced from the soft low noise audio coax (Belden 9239) to jacketed solder-soaked coax cable (Belden 1671J - RG405 compatible). The original cable indicated the delay of -34.3deg (@11MHz, 8.64ns) and the loss of 0.189dB.

I took 80-inch 1671J cable and measured the delay to be ~40deg. The length was adjusted using this number and the resulting cable indicated the delay of -34.0deg (@11MHz, 8.57ns) and the loss of 0.117dB.

The REFL11 demod module was restored and the cabling around REFL11 and AS110 were restored, tightened, and checked.

----

I've removed the PD mon cables from the NI RF switch. The open ports were plugged with 50Ohm temirnators.

----

I ask commissioners to make the final check of the REFL11 performance using CDS.

Attachment 1: IMG_0545.jpeg  708 kB  Uploaded Wed Mar 17 04:09:50 2021  | Hide | Hide all | Show all
IMG_0545.jpeg
Attachment 2: IMG_0547.jpeg  757 kB  Uploaded Wed Mar 17 04:10:04 2021  | Hide | Hide all | Show all
IMG_0547.jpeg
Attachment 3: D040179-A.pdf  278 kB  Uploaded Wed Mar 17 04:10:13 2021  | Show | Hide all | Show all
Attachment 4: IMG_0548.jpeg  459 kB  Uploaded Wed Mar 17 04:46:23 2021  | Hide | Hide all | Show all
IMG_0548.jpeg
Attachment 5: IMG_0550.jpeg  442 kB  Uploaded Wed Mar 17 04:46:23 2021  | Hide | Hide all | Show all
IMG_0550.jpeg
ELOG V3.1.3-